After finishing your first finance class, you decide to start your own investment plan. Your first stock to purchase is Walgreen’s (WAL). You are confident enough to project next year’s dividend as $1.00. However, you are unsure of future dividends, so you just assume they will grow at 5.00% per year going forward each year. If you want a 15.00% return on your investment, what value do you place on the stock today?
Answer #1

D1=Next Years Dividend =$1.00

g=Dividend growth rate =5%=0.05

R=Required return on investment =15%=0.15

Current Fair Price =P0

P0=D1/(R-g)

Current fair market price of the stock=1.00/(0.15-0.05)=$10.00

Value of the stock today=$10